The no. Of oxygen atoms present in 14.6gm of magnesium bicarbonate

Grade:11

2 Answers

mohan
7 years ago
HI FRIEND.......
Now iam solved it, I Got the answer is THREE
ThankS Friend...............
ALL THE BEST
yadnya yeolekar
6 years ago
molecular mass of magnesium bicarbonate = 146
no. of moles = given weight/moecular weight
                    =  14.6/146 =0.1                                                                              …...............(1)
        here, no. of moles of oxygen in magnesium bicarbonate = 16*3*2/16 = 6  …..............(2)
[from (1) and (2) ],
        0.1*6*NA =0.6 NA
 So, the number of oxygen atoms present in 14.6 g of magnesium bicarbonate is 0.6 NA.